A restaurant manager wants to purchase 200 sets of dishes. One design costs $25 per set, while the other costs $45 per set. If the manager has $7,400 to spend, how many of each design should be ordered?

Answer by stanbon(75887)

A restaurant manager wants to purchase 200 sets of dishes. One design costs $25 per set, while the other costs $45 per set. If the manager has $7,400 to spend, how many of each design should be ordered?
-----------------------------------
Let # of sets with 1st design be "a".
---
Let # of sets with 2nd design be "b".
---------------------------------------
Equations:
a + b = 200
25a + 45b = 7400
-------------------------
25a + 45(200-a) = 7400
25a + 45*200 - 45a = 7400
-20a = -1600
a = 80 sets of design # 1
----
Solve for "b":
b = 200-a
b = 200-80
b = 120 sets of design # 2
-----------------------------
Cheers,
Stan H.
===============